OpenStudy (geekfromthefutur):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (geekfromthefutur):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

firt 5 children... just put c=5 in the equation... 10a+6(5)=80 10a+30=80 10a-30+30=80-30 10a=50 10a/10=50/10 a=5 there are 5 children, 30 must be spent, 50 is left, since each adult costs 10, 5 there are 5 adult tickets

OpenStudy (yanasidlinskiy):

Got it? I seriously gotta go..

OpenStudy (geekfromthefutur):

ok ^-^ thx 4 da help

OpenStudy (yanasidlinskiy):

Glad I could help!:)